OpenStudy (anonymous):

Secretion of which hormone is NOT controlled by a negative-feedback mechanism? prolactin testosterone oxytocin

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

What do u think ?

OpenStudy (anonymous):

im thinkin c. for some reason

OpenStudy (abhisar):

May i know the reasons ?

OpenStudy (anonymous):

im guessing

OpenStudy (abhisar):

i think it should be prolactin. But i am not very sure

OpenStudy (anonymous):

no i got that one wrong but thanks for trying to help

OpenStudy (anonymous):

@Young_G_14 It is oxytocin.

OpenStudy (anonymous):

@Abhisar Oxytocin is the key hormone that drives childbrith and the separation of the placenta from the uterine wall. It is also involved in stimulating lactation.

OpenStudy (abhisar):

i know there is a positive feedback of oxytocin during lactation.

OpenStudy (anonymous):

dang i had a feeling it was c.

OpenStudy (abhisar):

@mrdoldum how is then prolactin controlled ?

OpenStudy (abhisar):

or regulated ?

OpenStudy (abhisar):

Oxytocin stimulates uterine contractions an initiates labor. this is called as foetal ejection reflex. But the point is how prolactin regulated then ?

OpenStudy (abhisar):

And if there is no negative feedback for oxytocin, then how it is regulated ?

OpenStudy (anonymous):

@abhisar I don't have my Vander's Human Physiology book with me (it is at work), but I believe that prolactin is kept "off" by certain hormones and that particular stimuli cause these regulatory hormones to be removed and allow prolactin to be produced. The negative part comes in, because I think (and I will get back to this question because I am not 100%), that as prolactin levels build up, the regulatory proteins begin to shut down production of prolactin. Now, for oxytocin, during childbirth the release of oxytocin causes more oxytocin to be released. This is pretty self evident by the increase in cervix dilation and other bodily functions for childbirth. A shot of oxytocin along with some other hormones is used to induce labor.

OpenStudy (anonymous):

Again, though, after a certain point other regulatory hormones block the transciption and translation of oxytocin. To be positive the release of the substance must stimulate increased production. Again, I am going to have to get back to you because I believe this question is requiring us to make a GROSS over-simplification.

OpenStudy (abhisar):

actually i read all these in first year...so need a brush up..i'll come back too after reading my guyton human physiology

OpenStudy (anonymous):

@Abhisar Found a pdf copy. So, prolactin is controlled via short-loop feedback and is inhibited by dopamine. It is what causes increased breast mass and milk production. However, the placenta causes the release of massive amounts of estrogen and progesterone which do not cause breast tissue growth to halt, but do stop milk production. Once the placenta is gone, milk production takes over. Stimulation of the nipple causes dopamine to be release to decrease and oxytocin sectrion. The dopamine causes the prolactin inhibotor to decreases, prolactin increases and milk synthesis occurs. At the same time oxytocin causes milk ejection form the mammary glands. Prolactin needs to remain largely unproduced except puberty, and reproduction so it is always inhibited, essentially.

OpenStudy (anonymous):

@Abhisar Now, on the side of oxytocin, its job is to start child birth and milk expulsion. During labor, mechanoreceptors in the uterine and vagina react to contractions causing the release of oxytocin that then increases the contractions. As for milk production, physical stimulation of the nipple causes oxytocin to be produced, and once again it causes more oxytocin production. In fact, women may even leak milk because hearing a baby cry can cause production and release of oxytocin.

OpenStudy (abhisar):

@mrdoldum actually feedback mechanism regulates the hormones under control of anterior pituitary. Oxytocin is secreted by posterior pituitary and is directly under neural control

OpenStudy (abhisar):

so yes the answer ton this question is oxytocin \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (anonymous):

@abhisar "These facts can now be put together in a unified pattern, as shown in Figure 19–28. Once started, the uterine contractions exert a positive-feedback effect upon themselves via both local facilitation of inherent uterine contractions and reflex stimulation of oxytocin secretion." Bit more complicated than absolute direct control in some cases.

OpenStudy (anonymous):

Vander's Human Phys. 8th edition.